Solve by substitution
x-y=3
3x-2y=3
Please show me if this problem is independent, dependent or inconsistent. More importantly I'd like to know the best way to solve future problems like this''.

Answer by ewatrrr(24785)


Hi,
x-y=3 OR y = x-3
3x - 2y = 3 |substituting for y
3x -2(x-3) = 3
3x - 2x + 6 = 3
x = 3 & y = 0 (x-3)
Line intersect at one point, Pt(3,0)
this is an Independent System of Equations